OpenStudy (anonymous):

Simplify. express in simplest radical form. (3-√2)(3+√2) thank you for all who help or suggest.

OpenStudy (jdoe0001):

recall that \(\bf (a-b)(a+b) = (a^2-b^2)\)

OpenStudy (cwrw238):

one way would be to use the identity (a - b)(a + b) = a^2 - b^2

OpenStudy (anonymous):

(3-2)(3+2)= (1)(5)=5?

OpenStudy (anonymous):

wait I did that wrong.

OpenStudy (cwrw238):

3^2 = 9 and (√2)^2 = ?

OpenStudy (anonymous):

(√2)^2=4

OpenStudy (cwrw238):

nope

OpenStudy (anonymous):

kinda solve like a linear equation...

OpenStudy (cwrw238):

just plug in a = 3 and √2 = b into RHS of the identity

OpenStudy (cwrw238):

and 2 * 2 = 4 not √2 * √2

OpenStudy (anonymous):

Sorry. Independent studies and I learn through visuals. have very bad reading comprehension skills

OpenStudy (cwrw238):

oh - i'm not sure how i can help you then

OpenStudy (anonymous):

not formilar with your RHS and what you mean by Identity when talking about this subject

OpenStudy (cwrw238):

oh sorry RHS is right hand side - in this case its a^2 - b^2 an identity is a mathematical statement which states one expression in an equivalent way and is true for all values of the variables a simple one would be 2(a + b) = 2a + 2b

OpenStudy (cwrw238):

so employing the identity to solve your problem we have (a - b)(a + b) = a^2 - b^2 and (3-√2)(3+√2) if we compare the 2 above we see that a = 3 and b = √2 so replacing a and b in RHS we get 3^2 - (√2)^2 and that simplifies to 9 - 2 = 7 and 7 is your simplest form

OpenStudy (anonymous):

you can totally be my tutor! your good. I understand now. I am a slow learner and I just don't have anyone in my family that could help me at this level of Algebra. So, I really appreciate the time you put in to help me better understand. thank you good sir.